### v1

# Generating Physically Stable and Buildable LEGO® Designs from Text

Abstract
We introduce LegoGPT, the first approach for generating physically stable LEGO brick models from text prompts. To achieve this, we construct a large-scale, physically stable dataset of LEGO designs, along with their associated captions, and train an autoregressive large language model to predict the next brick to add via next-token prediction. To improve the stability of the resulting designs, we employ an efficient validity check and physics-aware rollback during autoregressive inference, which prunes infeasible token predictions using physics laws and assembly constraints. Our experiments show that LegoGPT produces stable, diverse, and aesthetically pleasing LEGO designs that align closely with the input text prompts. We also develop a text-based LEGO texturing method to generate colored and textured designs. We show that our designs can be assembled manually by humans and automatically by robotic arms. We also release our new dataset, StableText2Lego, containing over 47,000 LEGO structures of over 28,000 unique 3D objects accompanied by detailed captions, along with our code and models at the project website: https://avalovelace1.github.io/LegoGPT/.
{strip}[Uncaptioned image]
Figure 1: Overview of LegoGPT. (a) Our method generates physically stable LEGO structures from text descriptions through an end-to-end approach, showing intermediate brick-by-brick steps. (b) The generated designs are buildable both by hand and by automated robotic assembly. (c) We show example results with corresponding text prompts. Besides basic LEGO designs (top), our method can generate colored LEGO models (bottom right) and textured models (bottom left) with appearance descriptions. We highly recommend the reader to check our website for step-by-step videos.
1 Introduction
3D generative models have made remarkable progress, driven by advances in generative modeling [15, 65] and neural rendering [26, 52]. These models have enabled various applications in virtual reality, gaming, entertainment, and scientific computing. For example, several works have explored synthesizing 3D objects from text [58], adding texture to meshes [61, 9], and manipulating the shape and appearance of existing 3D objects and scenes [40, 19].
However, applying existing methods to create real-world objects remains challenging. Most approaches focus on generating diverse 3D objects with high-fidelity geometry and appearance [85, 21], but these digital designs often cannot be physically realized due to two key challenges [46]. First, the objects may be difficult to assemble or fabricate using standard components. Second, the resulting structure may be physically unstable even if assembly is possible. Without proper support, parts of the design can collapse, float, or remain disconnected.
In this work, we address the challenge of generating physically realizable objects and study this problem in the context of LEGO® design. LEGO is widely used in entertainment, education, artistic creation, and manufacturing prototyping. Additionally, it can serve as a reproducible research benchmark, as all standard components are readily available. Due to the significant effort required for manual design, recent studies have developed automated algorithms to streamline the process and generate compelling results. However, existing approaches primarily create LEGO designs from a given 3D object [45] or focus on a single object category [13, 12].
Our goal is to develop a method for generating LEGO designs directly from freeform text prompts while ensuring physical stability and buildability. Specifically, we aim to train a generative model that produces designs that are:
  • Physically stable: Built on a LEGO baseplate with strong structural integrity, without floating or collapsing bricks.
  • Buildable: Compatible with standard LEGO pieces and able to be assembled brick-by-brick by humans or robots.
In this work, we introduce LegoGPT with the key insight of repurposing autoregressive large language models, originally trained for next-token prediction, for next-brick prediction. We formulate the problem of LEGO design as an autoregressive text generation task, where the next-brick dimension and placement are specified with a simple textual format. To ensure generated structures are both stable and buildable, we enforce physics-aware assembly constraints during both training and inference. During training, we construct a large-scale dataset of physically stable LEGO designs paired with captions. During autoregressive inference, we enforce feasibility with an efficient validity check and physics-aware rollback to ensure that the final tokens adhere to physics laws and assembly constraints.
Our experiments show that the generated designs are stable, diverse, and visually appealing while adhering to input text prompts. Our method outperforms pre-trained LLMs with and without in-context learning, and previous approaches based on mesh-based 3D generation. Finally, we explore applications such as text-driven LEGO texturing, as well as manual assembly and automated robotic assembly of our designs. Our dataset, code, and models are available at the project website: https://avalovelace1.github.io/LegoGPT/.
2 Related Work
Text-to-3D Generation.
Text-to-3D generation has seen remarkable progress in recent years, driven by advances in neural rendering and generative models. Dreamfusion [58] and Score Jacobian Chaining [76] pioneer zero-shot text-to-3D generation by optimizing neural radiance fields [52] with pre-trained diffusion models [62]. Subsequent work has explored alternative 3D representations [35, 3, 48, 66, 33, 42, 51] and improved loss functions [77, 83, 47, 74, 25, 44]. Rather than relying on iterative optimization, a promising alternative direction trains generative models directly on 3D asset datasets, with various backbones including diffusion models [63, 89, 32, 54, 85, 34, 20, 86, 60], large reconstruction models [21, 31, 81, 73], U-Nets [68, 36], and autoregressive models [55, 4, 67, 64, 5, 6, 18, 79].
However, these existing methods cannot be directly applied to LEGO generation because they do not account for the unique physical constraints and assembly requirements of real-world object designs [46]. Our work bridges this gap by introducing a framework for generating physically stable and buildable LEGO designs directly from text prompts.
Autoregressive 3D Modeling.
Recent research has successfully used autoregressive models to generate 3D meshes [55, 64, 4, 67, 5, 6, 18, 79], often conditioned on input text or images. Most recently, LLaMA-Mesh [78] demonstrates that large language models (LLMs) can be fine-tuned to output 3D shapes in plain-text format, given a text prompt. However, most existing autoregressive methods focus on mesh generation. In contrast, we focus on generating LEGO designs from text prompts, leveraging LLMs’ reasoning capabilities.
Figure 2: StableText2Lego Dataset. (a) From a ShapeNetCore [2] mesh, we generate a LEGO design by voxelizing it onto a $20\times 20\times 20$ grid and applying legolization to determine the brick layout. (b) We augment each shape with multiple structural variations by randomizing the brick layout while preserving the overall shape. (c) Stability analysis [37] is performed on each variation to filter out physically unstable designs. (d) To obtain the corresponding captions for each shape, we render the LEGO design from 24 different viewpoints and use GPT-4o [1] to generate detailed geometric descriptions. (e) Data samples from 5 categories in our StableText2Lego dataset.
LEGO Assembly and Generation.
Creating LEGO® designs given a reference 3D shape has been widely studied [28]. Existing works [57, 69, 88] formulate the generation as an optimization problem guided by hand-crafted heuristic rules. Such heuristics can include ensuring that all bricks are interconnected, minimizing the number of bricks, and maximizing the number of brick orientation alternations. Luo et al. [45] leverages structural stability estimation to find weak structural parts and rearrange the local brick layout to generate physically stable designs. Kim et al. [27], Liu et al. [39] formulate a planning problem to fill the target 3D model sequentially. However, these methods only generate designs given an input 3D shape, assuming a valid LEGO design exists, which is difficult to verify in practice.
Few works have explored learning-based techniques to generate LEGO designs. Thompson et al. [70] use a deep graph generative model in which the graph encodes brick connectivity. However, this method is limited to generating simple classes, such as walls and cuboids, using a single brick type. More recently, Ge et al. [13] use a diffusion model to predict a semantic volume, which is then translated into a high-quality micro building. Their method produces impressive results for a single category. Zhou et al. [87] and Ge et al. [12] generate compelling figurine designs given an input portrait. They use machine learning to select from a pre-made set of components that best match an input photo. Although effective for faces, extending this selection-based approach to arbitrary objects is challenging. Zhou et al. [90] formulate an optimization problem to create a LEGO model from an input image. While their output is a 2D LEGO mosaic, we focus on 3D structures in this work. Goldberg et al. [14] queries a vision-language model (VLM) to generate diverse 3D assembly structures. However, they use regular building blocks instead of LEGO bricks, which do not have interlocking connections and thus have limited expressiveness.
Our goal is closest to that of [30]. This work has three steps: (1) generating an image using a text-to-image model, (2) converting the image into voxels, and (3) using heuristics to create a physical LEGO build without considering physical constraints. In contrast, our method performs the text-to-LEGO task without requiring intermediate image or voxel representations.
Physics-Aware Generation.
Despite significant advances in 3D generation, incorporating physics constraints remains largely underexplored; most approaches focus primarily on geometry or appearance modeling. Physics-aware generation can be broadly categorized into two approaches: direct constraint enforcement and learned validation. Simple physical constraints, such as collision avoidance and contact requirements, can be incorporated directly through explicit penalty terms during optimization [53, 82, 84, 75, 41, 24, 16]. More complex physical properties, such as structural stability and dynamic behavior, typically require differentiable physics simulators [7, 56, 80, 50] or data-driven physics-aware assessment models [10, 49]. To our knowledge, our paper is the first attempt to incorporate physics-aware constraints into text-based LEGO generation.
3 Dataset
Training a modern autoregressive model requires a large-scale dataset. Therefore, we introduce StableText2Lego, a new large-scale LEGO® dataset that contains $47{,}000+$ different LEGO structures, covering $28{,}000+$ unique 3D objects from $21$ common object categories of the ShapeNetCore dataset [2]. We select categories featuring diverse and distinctive 3D objects while excluding those resembling cuboids. Each structure is paired with a group of text descriptions and a stability score, which indicates its physical stability and buildability. Below, we describe the dataset construction, an overview of which is given in Figure 2.
Figure 3: Method. (a) Our system tokenizes a LEGO design into a sequence of text tokens, ordered in a raster-scan manner from bottom to top. (b) We create an instruction dataset pairing brick sequences with descriptions to fine-tune LLaMA-3.2-Instruct-1B. (c) At inference time, LegoGPT generates LEGO designs incrementally by predicting one brick at a time given a text prompt. For each generated brick, we perform validity checks to ensure it is well-formatted, exists in our brick library, and does not collide with existing bricks. After completing the design, we verify its physical stability. If the structure is unstable, we roll back to the stable state by removing all the unstable bricks and their subsequent ones and resume generation from that point.
LEGO Representation.
We consider LEGO structures built on a LEGO baseplate. Each LEGO structure in StableText2Lego is represented as $B=[b_{1},b_{2},\dots,b_{N}]$ with $N$ bricks, and each element denotes a brick’s state as $b_{i}=[h_{i},w_{i},x_{i},y_{i},z_{i}]$. Here, $h_{i}$ and $w_{i}$ indicate the brick length in the $X$ and $Y$ directions, respectively, and $x_{i}$, $y_{i}$, and $z_{i}$ denote the position of the stud closest to the origin. The position has $x_{i}\in[0,1,\dots,H-1]$, $y_{i}\in[0,1,\dots,W-1]$, $z_{i}\in[0,1,\dots,D-1]$, and $H$, $W$, and $D$ represent the dimensions of the discretized grid world.
Mesh-to-LEGO.
We construct the dataset by converting 3D shapes from ShapeNetCore [2] into LEGO structures as shown in Figure 2(a). Given a 3D mesh, we voxelize and downsample it to a $20\times 20\times 20$ grid world to ensure a consistent scale, i.e., $H=W=D=20$. The LEGO brick layout is generated by a split-and-remerge legolization algorithm. To improve data quality and diversity, we introduce randomness in legolization and generate multiple different LEGO structures for the same 3D object, as illustrated in Figure 2(b). This increases the chance of obtaining a stable LEGO structure and more diverse LEGO brick layouts. More details can be found in Appendix A. We use eight commonly available standard bricks: $1\times 1$, $1\times 2$, $1\times 4$, $1\times 6$, $1\times 8$, $2\times 2$, $2\times 4$, and $2\times 6$.
Stability Score.
We assess the physical stability of each structure, as illustrated in Figure 2(c), using the analysis method [37]. For a structure $B=[b_{1},b_{2},\dots,b_{N}]$, the stability score $S\in\mathbb{R}^{N}$ assigns each brick $b_{i}$ a value $s_{i}\in[0,1]$ that quantifies the internal stress at its connections. Higher scores $s_{i}$ indicate greater stability, while $s_{i}=0$ indicates an unstable brick that will cause structural failure. Calculating the stability score requires solving a nonlinear program to determine the forces acting on each brick to achieve a static equilibrium that prevents structural collapse, as detailed in Section 4.2. For typically-sized (i.e., $<200$ bricks) structures in Figure 2, it takes $\sim$0.35 seconds on average. We only include stable structures where all bricks have stability scores greater than 0.
Caption Generation.
To obtain captions for each shape, we render the LEGO from 24 different viewpoints and combine them into a single multi-view image. We then prompt GPT-4o [1] to produce five descriptions for these renderings with various levels of detail. Importantly, we ask GPT-4o to omit color information and focus only on geometry. The complete GPT-4o prompt is provided in Appendix A.
Figure 2(e) shows several data samples in StableText2Lego. The rich variations within each category and the comprehensive text-LEGO pairs make it possible to train large-scale generative models. More insights on our StableText2Lego are discussed in Appendix A.
4 Method
Here, we introduce LegoGPT, a method for generating physically stable LEGO designs from text prompts. Leveraging LLMs’ ability to model sequences and understand text, we fine-tune a pre-trained LLM for the LEGO generation task (Section 4.1). To increase the stability and buildability of our designs, we use brick-by-brick rejection sampling and physics-aware rollback during inference (Section 4.2). Figure 3 illustrates an overview of our method.
4.1 Model Fine-tuning
Pre-trained LLMs excel at modeling sequences and understanding natural language, making them promising candidates for our task. We further fine-tune a pre-trained LLM on a custom instruction dataset containing text prompts and their corresponding LEGO structures from StableText2Lego.
Pre-trained Base Model.
We use LLaMA-3.2-1B-Instruct [11] as our base model. This model is fine-tuned to give coherent answers to instruction prompts, making it suitable for text-based LEGO design generation. As shown in Figure 5, the base model can generate LEGO-like designs through in-context learning, highlighting the promise of using pre-trained LLMs for our task. However, the generated designs often omit certain object parts and contain intersecting or disconnected bricks, making them physically unstable and unbuildable. To address these issues, we further fine-tune the pre-trained model using our dataset.
Instruction Fine-tuning Dataset.
For each stable design and its corresponding captions, we construct an instruction in the following format: “(user) Create a LEGO model of {caption}. (assistant) {lego-design}.”
To simplify training and reuse LLaMA’s tokenizer, we represent LEGO designs in plain text. But what format should we use? The standard LEGO design format LDraw [29] has two main drawbacks. First, it does not directly include brick dimensions, which are crucial for assessing the structure and validating brick placements. Second, it contains unnecessary information, such as brick orientation and scale. This information is redundant, as each axis-aligned brick has only two valid orientations.
[…]
With our fine-tuned model LegoGPT $\theta$, we predict the bricks $b_{1},b_{2},...,b_{N}$ in an autoregressive manner:
$p(b_{1},b_{2},...,b_{N}|\theta)=\prod_{i=1}^{N}p(b_{i}|b_{1},...,b_{i-1},\theta).$ (1)
4.2 Integrating Physical Stability
Although trained on physically stable data, our model sometimes generates designs that violate physics and assembly constraints. To address this issue, we further incorporate physical stability verification into autoregressive inference.
Figure 4: Force Model. (a) We consider all forces exerted on a single brick, including gravity (black), vertical forces with the top brick (red/blue) and bottom brick (green/purple), and horizontal (shear) forces due to knob connections (cyan), and adjacent bricks (yellow). (b) The structural force model $\mathcal{F}$ extends the individual force model to multiple bricks. Solving for static equilibrium in $\mathcal{F}$ determines each brick’s stability score.
A LEGO structure is considered physically stable and buildable if it does not collapse when built on a baseplate. To this end, we assess physical structural stability using the stability analysis method [37]. Here, we briefly overview this method below. Figure 4(a) illustrates all possible forces exerted on a single brick. We extend the single brick model and derive the structural force model $\mathcal{F}$, which consists of a set of candidate forces (e.g., pulling, pressing, supporting, dragging, normal, etc.), as demonstrated in Figure 4(b). For a LEGO structure $B=[b_{1},b_{2},\dots,b_{N}]$, each brick $b_{i}$ has $M_{i}$ candidate forces $F_{i}^{j}\in\mathcal{F}_{i},j\in[1,M_{i}]$. A structure is stable if all bricks can reach static equilibrium, i.e.,
$\sum_{j}^{M_{i}}F_{i}^{j}=0,\qquad\sum_{j}^{M_{i}}{\tau}_{i}^{j}\dot{=}\sum_{j}^{M_{i}}L_{i}^{j}\times F_{i}^{j}=0,$ (2)
where $L_{i}^{j}$ denotes the force lever corresponding to $F_{i}^{j}$. The stability analysis is formulated into a nonlinear program as
$\begin{split}\operatorname*{arg\,min}_{\mathcal{F}}&\sum_{i}^{N}\Biggl{\{}|\sum_{j}^{M_{i}}F_{i}^{j}|+|\sum_{j}^{M_{i}}{\tau}_{i}^{j}|+\alpha\mathcal{D}_{i}^{\max}+\beta\sum\mathcal{D}_{i}\Biggl{\}},\end{split}$ (3)
subject to three constraints: 1) all force candidates in $\mathcal{F}$ should take non-negative values; 2) certain forces exerted on the same brick cannot coexist, e.g., the pulling (red arrow) and pressing (blue arrow), the dragging (green arrow) and supporting (purple arrow); 3) Newton’s third law, e.g., at a given connection point, the supporting force on the upper brick should be equal to the pressing force on the bottom brick. $\mathcal{D}_{i}\subset\mathcal{F}_{i}$ is the set of candidate dragging forces (green arrow) on $b_{i}$. $\alpha$ and $\beta$ are hyperparameter weights.
Solving the above nonlinear program in Eqn. 3 using Gurobi [17] finds a force distribution $\mathcal{F}$ that drives the structure to static equilibrium with the minimum required internal stress, suppressing the overall friction (i.e., $\sum\mathcal{D}_{i}$) as well as avoiding extreme values (i.e., $\mathcal{D}_{i}^{\max}$). From the force distribution $\mathcal{F}$, we obtain the per-brick stability score as
$\begin{split}s_{i}=\begin{cases}0&\begin{array}[]{c}\sum_{j}^{M_{i}}F_{i}^{j}\neq 0\\ \vee\quad\sum_{j}^{M_{i}}{\tau}_{i}^{j}\neq 0\\ \vee\quad\mathcal{D}_{i}^{\max}>F_{T},\end{array}\\ \frac{F_{T}-\mathcal{D}_{i}^{\max}}{F_{T}}&\text{otherwise},\end{cases}\end{split}$ (4)
where $F_{T}$ is a measured constant friction capacity between brick connections. Higher scores $s_{i}$ indicate greater stability, while $s_{i}=0$ indicates an unstable brick that will cause structural failure: either $\mathcal{F}$ cannot reach static equilibrium ($\sum_{j}^{M_{i}}F_{i}^{j}\neq 0\vee\sum_{j}^{M_{i}}\tau_{i}^{j}\neq 0$) or the required friction exceeds the friction capacity of the material ($\mathcal{D}_{i}^{\max}>F_{T}$). Due to the equality constraints imposed by Newton’s third law, Eqn. 3 includes only the dragging forces and excludes pulling forces. For a physically stable structure, we need $s_{i}>0,\forall i\in[1,N]$.
When to apply stability analysis?
Our model generates LEGO designs sequentially, one LEGO brick at a time. A straightforward approach to ensuring physical stability is to apply stability analysis to each step and resample a brick that would cause a collapse. However, this step-by-step validation, though efficient per check, could be time-consuming due to the large number of checks required. More importantly, many LEGO designs are unstable when partially constructed, yet become stable when fully assembled. Adding a stability check after each brick generation could overly constrain the model exploration space. Instead, we propose brick-by-brick rejection sampling combined with physics-aware rollback to balance stability and diversity.
Brick-by-Brick Rejection Sampling
To improve inference speed and avoid overly constraining the model generation, we relax our constraints during inference. First, when the model generates a LEGO brick and its position, the brick should be well-formatted (e.g., available in the inventory) and not lie outside the workspace. Second, we ensure that newly added bricks do not collide with the existing structure. Formally, for each generated brick $b_{t}$, we have $\mathcal{V}_{t}\cap\mathcal{V}_{i}=\varnothing,\forall i\in[1,t-1]$, where $\mathcal{V}_{i}$ denotes the voxels occupied by $b_{i}$. These heuristics allow us to efficiently generate well-formatted LEGO structures without explicitly considering complex physical stability. To integrate these heuristics, we use rejection sampling: if a brick violates the heuristics, we resample a new brick from the model. Due to the relaxed constraints, most bricks are valid, and rejection sampling does not significantly affect inference time.
Physics-Aware Rollback
To ensure that the final design $B=[b_{1},b_{2},\dots,b_{N}]$ is physically stable, we calculate the stability score $S$. If the resulting design is unstable, i.e., $s_{i}=0,i\in\mathcal{I}$, we roll back the design to the state before the first unstable brick was generated, i.e., $B^{\prime}=[b_{1},b_{2},\dots,b_{\min{\mathcal{I}}-1}]$. Here, $\mathcal{I}$ is the set of the indices of all the unstable bricks. We repeat this process iteratively until we reach a stable structure $B^{\prime}$, and continue generation from the partial structure $B^{\prime}$. Note that we can use the per-brick stability score to efficiently find the collapsing bricks and their corresponding indices in the sequence. We summarize our inference sampling in Algorithm 1.
Input: Text prompt $c$; Autoregressive model $\theta$
Output: LEGO design following the text prompt
1 $B\leftarrow\text{empty LEGO structure}$;
2 loop
       /* Predict next brick w/ rejection sampling */
[…]
6             if $b$ is valid then break;
7
8       end for
9      $B.\texttt{add_brick}(b)$;
10       if $b$ contains EOF then // Structure complete
11             if $B$ is stable or max rollbacks exceeded then return $B$;
12             while $B$ is unstable do // Rollback if unstable
13                   $\mathcal{I}\leftarrow$ indices of unstable bricks in $B$;
                   $i\leftarrow\min\mathcal{I}$;
                   // idx of 1st unstable brick
14                   $B\leftarrow[b_{1},\dots,b_{i-1}]$;
15
16             end while
17
18       end if
19
20 end
21
ALGORITHM 1 LegoGPT inference algorithm.
Table 1: Quantitative Analysis. We evaluate our method against several baselines on validity (no out-of-library, out-of-bounds, or colliding bricks), stability, and CLIP-based text similarity. Stability and CLIP score are computed over valid designs only. For LLaMA-Mesh [78], validity requires a well-formed OBJ file. Our method outperforms all baselines as well as the ablated setups on validity and stability using our proposed rejection sampling and rollback, while maintaining high text similarity.
Method % valid $\uparrow$ % stable $\uparrow$ mean brick stability $\uparrow$ min brick stability $\uparrow$ CLIP $\uparrow$
Pre-trained LLaMA (0-shot) 0.0% 0.0% N/A N/A N/A
In-context learning (5-shot) 2.4% 1.2% 0.675 0.479 0.284
LLaMA-Mesh [78] 94.8% 50.8% 0.894 0.499 0.311
LGM [68] 100% 25.2% 0.942 0.231 0.299
XCube [60] 100% 75.2% 0.964 0.686 0.322
Hunyuan3D-2 [86] 100% 75.2% 0.973 0.704 0.324
Ours w/o rejection sampling or rollback 37.2% 12.8% 0.956 0.325 0.329
Ours w/o rollback 100% 24.0% 0.947 0.228 0.322
Ours (LegoGPT) 100% 98.8% 0.996 0.915 0.324
Figure 5: Result gallery and baseline comparisons. Our method can generate high-quality, diverse, and novel LEGO designs aligned with the given text prompts. Black bricks are colliding. For LLaMA-Mesh [78], LGM [68], XCube [60], and Hunyuan3D-2 [86], an inset of the generated mesh before legolization is shown.
4.3 LEGO Texturing and Coloring
While our primary focus is on generating LEGO shapes, color and texture play a critical role in creative LEGO designs. Therefore, we propose a method that applies detailed UV textures or assigns uniform colors to individual bricks.
UV Texture Generation.
Given a structure $B$ and its corresponding mesh $\mathcal{M}$, we first identify the set of occluded bricks $B_{\text{occ}}$ that have all six faces covered by adjacent bricks, and remove $B_{\text{occ}}$ for efficiency. The remaining bricks $B_{\text{vis}}=B\setminus B_{\text{occ}}$ are merged into a single mesh $\mathcal{M}$ with cleaned overlapping vertices using ImportLDraw [72]. We generate a UV map $\text{UV}_{\mathcal{M}}$ by cube projection. The texture map $I_{\text{texture}}$ is then generated using FlashTex [9], a fast text-based mesh texturing approach:
$I_{\text{texture}}=\text{FlashTex}(\mathcal{M},\text{UV}_{\mathcal{M}},c),$ (5)
where text prompt $c$ describes the visual appearance. This texture can be applied through UV printing or stickers.
Uniform Brick Color Assignment.
We can also assign each brick a uniform color from the LEGO color library. Given a structure $B$, we convert it to a voxel grid $\mathcal{V}$ and then to a UV-unwrapped mesh $\mathcal{M}_{\mathcal{V}}$. For every voxel $v\in\mathcal{V}$, let $f^{v}_{i},i=1,\dots,N_{v}$ be its visible faces where $0\leq N_{v}\leq 6$. Each face $f^{v}_{i}$ is split into two triangles and mapped to a UV region $\mathcal{S}^{v}_{i}$, creating a mesh $\mathcal{M}_{\mathcal{V}}$ with UV map $\text{UV}_{\mathcal{V}}$. We apply FlashTex [9] to generate a texture $I_{\text{texture}}$:
$I_{\text{texture}}=\text{FlashTex}(\mathcal{M}_{\mathcal{V}},\text{UV}_{\mathcal{V}},c).$ (6)
The color of each voxel $\mathcal{C}(v)\in\mathbb{R}^{3}$ is computed as:
$\mathcal{C}(v)=\frac{1}{N_{v}}\sum_{i=1}^{N_{v}}\mathcal{C}(f^{v}_{i}),\quad\forall v\in\mathcal{V},$ (7)
where $\mathcal{C}(f^{v}_{i})=\frac{1}{|\mathcal{S}^{v}_{i}|}\sum_{(x,y)\in\mathcal{S}^{v}_{i}}I_{\text{texture}}(x,y)$ is the color of each visible face $f^{v}_{i}$, and $|\mathcal{S}^{v}_{i}|$ represents the number of pixels in region $\mathcal{S}^{v}_{i}$ in the UV map. For each brick $b_{t}$ and its constituent voxels $\mathcal{V}_{t}$, we compute the brick color $\mathcal{C}(b_{t})=\frac{1}{|\mathcal{V}_{t}|}\sum_{v\in\mathcal{V}_{t}}\mathcal{C}(v)$. Finally, we find the closest color in the color set. While UV texturing offers higher-fidelity details, uniform coloring allows us to use standard LEGO bricks.
Figure 6: Ablation study. Brick-by-brick rejection sampling and physics-informed rollback help to ensure that the generated LEGO design is both valid and stable. Black indicates colliding bricks.
Figure 7: Textured and Colored LEGO Generation. Our method can generate diverse textured (top two rows) and colored (bottom) LEGO models based on the same shape while using different appearance text prompts.
Figure 8: Automated Assembly. Demonstrations of automated assembly of generated LEGO structures using robots.
5 Experiments
5.1 Implementation Details
Fine-tuning.
Our fine-tuning dataset contains 240k distinct prompts and 47k+ distinct LEGO designs. We use 90% of the data for training and 10% for evaluation. For efficiency, we include samples only up to 4096 tokens in length. Training details are provided in Appendix A.
Inference.
To evaluate our method, we generate one LEGO design for each of 250 prompts randomly selected from the validation dataset. The nonlinear optimization in Eqn. 3 is solved using Gurobi [17]. We set $F_{T}=0.98N$ with $\alpha=10^{-3}$ and $\beta=10^{-6}$. We allow up to 100 physics-aware rollbacks before accepting the LEGO design. The median number of required rollbacks is 2, and the median time to generate one LEGO design is 40.8 seconds.
5.2 LEGO Generation Results
Figure 5 shows a gallery of diverse, high-quality LEGO designs that closely follow the input prompts.
Baseline Comparisons.
As baselines, we use LLaMA-Mesh [78], LGM [68], XCube [60], and Hunyuan3D-2 [86] to generate a mesh from each prompt, and then convert the meshes to LEGO via legolization. Additionally, we compare our method with pre-trained models evaluated in both a zero-shot and few-shot manner. For few-shot evaluation, we provide the model with 5 examples of stable LEGO designs and their captions. We cannot run comparisons with Lennon et al. [30] as the code has not been released.
For each method, we compute the proportion of stable and valid structures among the generated designs. Additionally, for each valid structure, we compute its mean and minimum brick stability scores. To evaluate prompt alignment, we compute the CLIP score [59] between a rendered image of each valid structure and the text “A LEGO model of {prompt}”. As shown in Table 1, our method outperforms all baselines in these metrics. Figure 5 shows that our method generates LEGO structures of higher quality than the baselines.
Ablation Study.
We demonstrate the importance of rejection sampling and physics-aware rollback. As seen in Figure 6, rejection sampling eliminates invalid bricks, such as those with collisions, while rollback helps to ensure the final model is physically stable. The quantitative results in Table 1 show that our full method generates a higher proportion of valid and stable LEGO designs, while closely following the text prompts.
5.3 Extensions and Applications
Robotic Assembly of Generated LEGO.
We demonstrate automated assembly using a dual-robot-arm system in Figure 8. The robots use the manipulation policy [38] and the asynchronous multi-agent planner [23] to manipulate LEGO bricks and construct the structure. More details are included in Appendix B. Since the generated structures are physically stable, efficient and automated assembly can be performed.
Manual Assembly.
Our generated LEGO structures are physically valid and can be assembled by hand. See Appendix C for details.
Textured and Colored LEGO Model Generation.
Figure 7 shows both UV texturing and uniform coloring results of LEGO models, demonstrating our method’s ability to generate diverse styles while preserving the underlying geometry.
6 Discussion and Limitations
In this work, we have introduced LegoGPT, an autoregressive model for generating LEGO designs from text prompts. Our method learns to predict the next brick sequentially while ensuring physical stability and buildability. We have shown that our method outperforms LLM backbone models and several recent text-to-3D generation methods.
Limitations.
Though our method outperforms existing methods, it still has several limitations. First, due to limited computational resources, we have not explored the largest 3D dataset. As a result, our method is restricted to generating designs within a $20\times 20\times 20$ grid across 21 categories. Future work includes scaling up model training on larger, more diverse datasets, such as Objaverse-XL [8] at higher grid resolutions. Training on large-scale datasets can also improve generalization to out-of-distribution text prompts.
Second, our method currently supports a fixed set of commonly used LEGO bricks. In future work, we plan to expand the brick library to include a broader range of dimensions and brick types, such as slopes and tiles, allowing for more diverse and intricate LEGO designs.
Acknowledgments.
We thank Minchen Li, Ken Goldberg, Nupur Kumari, Ruihan Gao, and Yihao Shi for their discussions and help. We also thank Jiaoyang Li, Philip Huang, and Shobhit Aggarwal for developing the bimanual robotic system. This work is partly supported by the Packard Foundation, Cisco Research Grant, and Amazon Faculty Award. This work is also in part supported by the Manufacturing Futures Institute, Carnegie Mellon University, through a grant from the Richard King Mellon Foundation. KD is supported by the Microsoft Research PhD Fellowship.
Appendix
Appendix A LegoGPT Implementation Details
Captioning Details.
The complete prompt template used for GPT-4o caption generation is as follows:
“This is a rendering of a 3D object built with LEGO bricks with 24 different views. The object belongs to the category of {CATEGORY_NAME}. You will generate five different captions for this {CATEGORY_NAME} that:
1. Describes the core object/subject and its key geometric features
2. Focuses on structure, geometry, and layout information
3. Uses confident, concrete, and declarative language
4. Omits color and texture information
5. Excludes medium-related terms (model, render, design)
6. Do not describe or reference each view individually.
7. Focus on form over function. Describe the physical appearance of components rather than their purpose.
8. Describe components in detail, including size, shape, and position relative to other components.
9. The five captions should be from coarse to fine, with the first one being the most coarse-grained (e.g., a general description of the object, within 10 words) and the last one being the most fine-grained (e.g., a detailed description of the object, within 50 words). The five captions should be different from each other. Do not include any ordering numbers (e.g., 1, a, etc.).
10. Describe the object using the category name "{CATEGORY_NAME}" or synonyms of the category name "{CATEGORY_NAME}".”
StableText2Lego Details.
We generate StableText2Lego from the objects in ShapeNetCore [2]. While ShapeNetCore provides both voxel and mesh representations, we find that working directly with mesh data preserves better geometric details. We voxelize the 3D mesh into a $20\times 20\times 20$ grid representation and generate the LEGO structure using a split-and-remerge legolization algorithm. Split-and-remerge legolization has been studied in prior work [45]. But instead of initializing the voxel with $1\times 1$ unit bricks and randomly merging them into larger bricks [45], we initialize with heuristics by directly placing bricks to fill all the voxels, i.e., prioritizing placing 1) larger bricks and 2) bricks that connect multiple other bricks. After initializing the layout, we apply the new stability analysis [37] to identify weak parts in the structure and rearrange the local brick layout if necessary. In addition, to further improve data quality and diversity, we introduce randomness in the legolization process. Specifically, when initializing the LEGO structure, we randomly permute the order of brick placement within the same heuristic priority and generate two different structures for each object. With the mesh-to-LEGO and structure augmentation techniques, we generate $62{,}000+$ LEGO structures covering the $21$ categories with $31{,}218$ unique 3D objects from ShapeNetCore. Among the selected 3D objects, $\sim$92(i.e., $28{,}822$) of them have at least one stable LEGO design, which offer $47{,}000+$ stable LEGO layouts. Our StableText2Lego dataset significantly expands upon the previous StableLego dataset [37] in several key aspects: it contains >3×>3$\times$> 3 × more stable unique 3D objects with >5×>5$\times$> 5 × more stable LEGO structures compared to $8{,}000+$ in StableLego, spans a diverse set of 21 object categories, and provides detailed geometric descriptions for each shape.
Figure 9: Given a prompt and several example LEGO designs in text format, a pre-trained LLaMA model can generate LEGO designs with some structure.
Figure 10: Novelty analysis. For each structure generated by LegoGPT, we find the closest structure in the training dataset as measured by Chamfer distance in voxel space. The generated structures are distinct from their nearest neighbors, indicating low memorization.
Training details.
The full prompt that we use to construct our instruction fine-tuning dataset is as follows:
[SYSTEM]You are a helpful assistant.
[USER]Create a LEGO model of the input. Format your response as a list of bricks: <brick dimensions> <brick position>, where the brick position is (x,y,z).
Allowed brick dimensions are 2x4, 4x2, 2x6, 6x2, 1x2, 2x1, 1x4, 4x1, 1x6, 6x1, 1x8, 8x1, 1x1, 2x2. All bricks are 1 unit tall.
### Input:
[INSERT CAPTION]
We fine-tune using low-rank adaptation (LoRA) [22] with a rank of 32, alpha of 16, and dropout rate of 0.05. To prevent catastrophic forgetting and training instability, we apply LoRA to only the query and value matrices for a total of 3.4M tunable parameters. We train for three epochs on eight NVIDIA RTX A6000 GPUs. We use the AdamW optimizer [43] with a learning rate of $0.002$, using a cosine scheduler with 100 warmup steps and a global batch size of 64. The total training time is 12 hours.
Inference Details.
The sampling temperature is set to 0.6 for all experiments. During brick-by-brick rejection sampling, to mitigate repeated generation of rejected bricks, we increase the temperature by 0.01 each time the model generates a brick that has already been rejected.
To force each output brick to be in the format “{h}×{w} ({x},{y},{z})”, we sample only from the set of valid tokens during each step. For example, the first token must be a digit, the second an ‘x’, and so on. This has little to no effect on the output of LegoGPT, but helps force the baseline pre-trained models to output LEGO bricks in the specified format.
Only 1.2% of the generated designs exceeded the maximum number of rollbacks and failed to produce a stable final structure.
Novelty Analysis.
For each generated structure, we find its closest structure in the training dataset, measured by computing the Chamfer distance in voxel space. As seen in Figure 10, the generated structures are distinct from their nearest neighbors in the dataset, confirming that our model can create novel designs rather than simply memorizing the training data.
In-context Learning.
We use LLaMA-3.2-1B-Instruct [11] as our base model, chosen for its computational efficiency. The in-context learning pipeline is shown in Figure 9, where the base model can generate LEGO-like designs through in-context learning, while suffering from collisions and disconnectivity. We did not use rejection sampling or rollback when evaluating zero-shot or few-shot generation, as doing so resulted in an excessive number of rejections and a sharp increase in generation time.
Appendix B Robotic Assembly
We demonstrate automated assembly using a dual-robot-arm system as shown in Figure 8. The system consists of two Yaskawa GP4 robots, each equipped with an ATI force-torque sensor. A calibrated LEGO plate is placed between them, and the robots use the bricks initially placed on the plate to construct the LEGO structure. Given a $B=[b_{1},b_{2},\dots,b_{N}]$, we employ the action mask in [39] with assembly-by-disassembly search [71] to generate a physically executable assembly sequence for the robots, i.e., reordering the brick sequence so that 1) each intermediate structure is physically stable by itself, 2) each intermediate structure is stable under the robot operation, and 3) each assembly step is executable within the system’s dexterity. With the executable assembly sequence, an asynchronous planner [23] distributes the assembly tasks to the robots, plans the robots’ movements, and coordinates the bimanual system to construct the LEGO structure. The robots use the end-of-arm tool and the manipulation policy presented in [38, 23] with closed-loop force control to robustly manipulate LEGO bricks and construct the structure.
Appendix C Manual Assembly
As shown in Figure 11, human users can assemble our generated structures, demonstrating their physical stability and buildability. Notably, since our method outputs a sequence of intermediate steps, it naturally serves as an intuitive assembly guide.
Figure 11: Manual Assembly. Demonstrations of manual assembly of generated LEGO structures.